OpenStudy (anonymous):

7a-b=15a+12, for a please show work

OpenStudy (mayankdevnani):

find value??

OpenStudy (anonymous):

yes for a

OpenStudy (anonymous):

Subtract 15a both the sides first and tell me what do you get ??

OpenStudy (mayankdevnani):

7a-b-15a=12 -8a-b=12 -8a=12+b -a=12+b/8 a=-12+b/8

OpenStudy (mayankdevnani):

ok @lyserdoll

OpenStudy (anonymous):

Stop doing this @mayankdevnani

OpenStudy (anonymous):

And you are not writing answer in proper form too.. You are not using any parenthesis: According to you : \[a = -12 + \frac{b}{8}\] That is totally wrong..

OpenStudy (mayankdevnani):

why?

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

so b=-8a-12

OpenStudy (mayankdevnani):

then we find the value of a

OpenStudy (anonymous):

yaa??? You are supposed to write your answer. if you have given directly, in proper form using brackets so that the asker does not understand it in a wrong way...

OpenStudy (mayankdevnani):

ok sorry

OpenStudy (mayankdevnani):

so what's the answer

OpenStudy (anonymous):

yeah im really confused right now

OpenStudy (anonymous):

Don't be sorry... Just give the right solution to the asker even if you are giving answer in one go..

OpenStudy (mayankdevnani):

@waterineyes what's the answer

OpenStudy (anonymous):

\[a = - \frac{(12+b)}{8}\]

OpenStudy (anonymous):

Isn't it ???

OpenStudy (anonymous):

Gotta go..

OpenStudy (mayankdevnani):

yaa m also

OpenStudy (mayankdevnani):

in short m not using brackets

OpenStudy (mayankdevnani):

it can also be a=\[\frac{-b}{8}-\frac{3}{2}\]

OpenStudy (anonymous):

@Razzputin stfu he can do what he wanna do leave him alone